# Stacked Pseudo-Convergent Sequences and Polynomial Dedekind Domains

### v2

@@ 2.3 Residually algebraic torsion extensions of $\mathbb{Z}_{(p)}$
The next proposition is analogous to Proposition 2.4, and characterizes residually algebraic torsion extensions of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$ in terms of pseudo-convergent sequences of $\overline{\mathbb{Q}_{p}}$ which are of transcendental type over $\mathbb{Q}$; clearly, every pseudo-convergent sequence of transcendental type of $\overline{\mathbb{Q}_{p}}$ belongs to this class. As a particular case, we find again part of the result of [21, Theorem 2.5].
@@
If $\widehat{\mathbb{Q}(X)}$ is a finite extension of $\mathbb{Q}_{p}$ then clearly $X$ is algebraic over $\mathbb{Q}_{p}$, so 1. implies 2. If $X$ is algebraic over $\mathbb{Q}_{p}$, we may identify $X$ with some $\alpha\in\overline{\mathbb{Q}_{p}}$; $\mathbb{Q}_{p}(\alpha)$ is a finite extension of $\mathbb{Q}_{p}$, hence complete. So, $\widehat{\mathbb{Q}(X)}=\mathbb{Q}_{p}(\alpha)$. As in the proof of [21, Theorem 2.5] it follows easily that $W=\mathbb{Z}_{(p),\alpha}$. Therefore, 2. implies 3.
If $W=\mathbb{Z}_{(p),\alpha}$ for some $\alpha\in\overline{\mathbb{Q}_{p}}$ transcendental over $\mathbb{Q}$, then by [21, Proposition 2.2], $ef<\infty$, so 3. implies 4. Finally, 4. implies 3. by [21, Lemma 2.4].
Note that if $E\subset\overline{\mathbb{Q}_{p}}$ is a pseudo-convergent sequence converging to $\alpha$ (hence, $E$ is a Cauchy sequence, which is of transcendental type over $\mathbb{Q}$ but of algebraic type over $\mathbb{Q}_{p}$), then $W=\mathbb{Z}_{(p),E}=\mathbb{Z}_{(p),\alpha}$ (see §1.1).
The claims about the value group and residue field of $\mathbb{Z}_{(p),\alpha}$ follow by [21, Proposition 2.2].
@@
Let $W$ be a residually algebraic torsion extension of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$. Then the completion $\widehat{\mathbb{Q}(X)}$ with respect to $W$ is a subfield of $\mathbb{C}_{p}$ if and only if there exists $\alpha\in\mathbb{C}_{p}$, transcendental over $\mathbb{Q}$, such that $W=\mathbb{Z}_{(p),\alpha}$.
@@
According to Proposition 2.24, when passing to the completion, either $X$ is algebraic over $\mathbb{Q}_{p}$ or $X$ is transcendental over $\mathbb{Q}_{p}$ and consequently either $\widehat{\mathbb{Q}(X)}\subset\overline{\mathbb{Q}_{p}}\subset\mathbb{C}_{p}$ or $\mathbb{Q}_{p}(X)\subset\widehat{\mathbb{Q}(X)}$. In the first case, $W=\mathbb{Z}_{(p),\alpha}$ for some $\alpha\in\overline{\mathbb{Q}_{p}}\subset\mathbb{C}_{p}$ transcendental over $\mathbb{Q}$. In the second case, $\widehat{\mathbb{Q}_{p}(X)}=\widehat{\mathbb{Q}(X)}$, where the completion of $\mathbb{Q}_{p}(X)$ is considered with respect to the valuation domain $\widetilde{W}=\widehat{W}\cap\mathbb{Q}_{p}(X)$. In particular, by Proposition 2.23, we get that $\widehat{\mathbb{Q}(X)}\subseteq\mathbb{C}_{p}$ if and only if there exists a transcendental element $\alpha\in\mathbb{C}_{p}$ such that $W=\mathbb{Z}_{(p),\alpha}$. ∎
@@
The following result is the analogous of Theorem 2.21 for building residually algebraic torsion extensions $W$ of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$ with prescribed residue field $k$ and value group $\Gamma$. Note that, contrary to that Theorem, now we are not assuming anymore that $[k:\mathbb{F}_{p}]\cdot[\Gamma:\mathbb{Z}]=\infty$.
@@
Let $e=[\Gamma:\mathbb{Z}]$ and $f=[k:\mathbb{F}_{p}]$. If $ef<\infty$, then it is well known that there exists $\alpha\in\overline{\mathbb{Q}_{p}}$ transcendental over $\mathbb{Q}$ such that $\mathbb{O}_{p,\alpha}$ has residue field $k$ and value group $\Gamma$. Hence, by [21, Proposition 2.2], $\mathbb{Z}_{(p),\alpha}$ is the desired extension of $\mathbb{Z}_{(p)}$.
@@
Therefore, if we choose $a_{i}\in\mathbb{Q}$ sufficiently $v_{p}$-adically close to $\alpha_{i}$ for each $i=0,\ldots,d$, we have $w(f-g)\geq 0$. In particular, $g\in W\cap\mathbb{Q}(X)=\mathbb{Z}_{(p),\alpha}$. The polynomial $h=f-g$ is in $\mathbb{Z}_{p}[X]$; therefore $f(\alpha)=h(\alpha)+g(\alpha)\in\mathbb{O}_{p}$, so that $f\in\mathbb{Z}_{p,\alpha}$. Therefore $W\cap\mathbb{Q}_{p}[X]\subseteq\mathbb{Z}_{p,\alpha}\cap\mathbb{Q}_{p}[X]$. Similarly, one can easily show that the other containment holds, so $W\cap\mathbb{Q}_{p}[X]=\mathbb{Z}_{p,\alpha}\cap\mathbb{Q}_{p}[X]$. In the same way, $M_{W}\cap\mathbb{Q}_{p}[X]=M_{p,\alpha}\cap\mathbb{Q}_{p}[X]$, where $M_{W}$ and $M_{p,\alpha}$ are the maximal ideals of $W$ and $\mathbb{Z}_{p,\alpha}$, respectively.
@@
Let $f=[W/M:\mathbb{Z}/p\mathbb{Z}]$. If $f<\infty$, then the claim follows by [21, Theorem 2.5] and corresponds to the first case of Proposition 2.24: $W=\mathbb{Z}_{(p),\alpha}$, for some $\alpha\in\overline{\mathbb{Q}_{p}}$ which is transcendental over $\mathbb{Q}$. If $f=\infty$, then we are in the last case of Proposition 2.24, so $W=\mathbb{Z}_{(p),E}$, for some pseudo-convergent sequence in $\overline{\mathbb{Q}_{p}}$ of transcendental type. As in the proof of Proposition 2.24, we denote by $\widehat{W}$ the completion of $W$; since the ramification index $e(W\mid\mathbb{Z}_{(p)})$ is finite, $\widetilde{W}=\widehat{W}\cap\mathbb{Q}_{p}(X)$ is a residually algebraic torsion extension of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$ which is a DVR, so by Theorem 2.17, $\widetilde{W}=\mathbb{Z}_{p,\alpha}$ for some $\alpha\in\mathbb{C}_{p}^{\text{br}}\setminus\overline{\mathbb{Q}_{p}}$. Hence, $W=\widetilde{W}\cap\mathbb{Q}(X)=\mathbb{Z}_{(p),\alpha}$. Note that $\alpha$ is transcendental over $\mathbb{Q}_{p}$, hence also over $\mathbb{Q}$.
@@
Conversely, suppose that $\mathbb{Z}_{(p),\alpha}=\mathbb{Z}_{(p),\beta}$. By Lemma 2.27, it follows that $\mathbb{Z}_{p,\alpha}=\mathbb{Z}_{p,\beta}$. Note that the last two valuation domains are the contraction to $\mathbb{Q}_{p}(X)$ of the valuation domains $\overline{\mathbb{Z}_{p}}_{\alpha}=\{\phi\in\overline{\mathbb{Q}_{p}}(X)\mid\phi(\alpha)\in\mathbb{O}_{p}\},\overline{\mathbb{Z}_{p}}_{\beta}=\{\phi\in\overline{\mathbb{Q}_{p}}(X)\mid\phi(\alpha)\in\mathbb{O}_{p}\}$ of $\overline{\mathbb{Q}_{p}}(X)$, respectively. By [6, 4, Chapt. VI, §8, 6., Corollary 1], there exists a $\mathbb{Q}_{p}(X)$-automorphism $\sigma$ of $\overline{\mathbb{Q}_{p}}(X)$ such that $\sigma(\overline{\mathbb{Z}_{p}}_{\alpha})=\overline{\mathbb{Z}_{p}}_{\beta}$. It is easy to check that $\sigma(\overline{\mathbb{Z}_{p}}_{\alpha})=\overline{\mathbb{Z}_{p}}_{\sigma(\alpha)}$. In particular, $\overline{\mathbb{Z}_{p}}_{\sigma(\alpha)}=\overline{\mathbb{Z}_{p}}_{\beta}$. If $\sigma(\alpha)-\beta\not=0$, let $c\in\overline{\mathbb{Z}_{p}}$ be such that $v_{p}(c)>v_{p}(\sigma(\alpha)-\beta)$. Let $a\in\overline{\mathbb{Q}_{p}}$ be such that $v_{p}(a-\sigma(\alpha))\geq v_{p}(c)$. Then the polynomial $\frac{X-a}{c}$ is in $\overline{\mathbb{Z}_{p}}_{\sigma(\alpha)}$ and not in $\overline{\mathbb{Z}_{p}}_{\beta}$, a contradiction. ∎
@@ 3 Polynomial Dedekind domains
We remark that we have the following representation for the ring $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$ as an intersection of valuation overrings (see [19, (2.2)], for example):
@@
where $\mathcal{P}^{\text{irr}}$ denotes the set of irreducible polynomials in $\mathbb{Q}[X]$. By [21, Proposition 2.2], the valuation domain $\mathbb{Z}_{(p),\alpha_{p}}$ of $\mathbb{Q}(X)$ has rank one if and only if $\alpha_{p}$ is transcendental over $\mathbb{Q}$, and has rank $2$ otherwise (in the last case, note that necessarily $\alpha\in\overline{\mathbb{Q}_{p}}$).
A totally similar argument to [19, Lemma 2.5] shows that, for $p\in\mathbb{P}$, we have
@@
Suppose that the above conditions on $\underline{E}$ are satisfied. By (3.1), $R$ is equal to an intersection of DVRs. Moreover, $R$ has finite character, that is, for every non-zero $f\in R$, $f$ belongs to finitely many maximal ideals of the family of DVRs appearing in (3.1): in fact, if $f(X)=\frac{g(X)}{n}$, for some $g\in\mathbb{Z}[X]$ and $n\in\mathbb{Z}$, $n\not=0$, then $f$ is divisible only by finitely many $q\in\mathcal{P}^{\text{irr}}$; since $\underline{E}$ is polynomially factorizable, by [19, Lemma 2.12], the set $\{p\in\mathbb{P}\mid\exists\alpha_{p}\in E_{p},v_{p}(g(\alpha_{p}))>0\}$ is finite, so that $f$ belongs to finitely many maximal ideals of the family $\mathbb{Z}_{(p),\alpha_{p}}$, $\alpha_{p}\in E_{p},p\in\mathbb{P}$. Hence, $R$ is a Krull domain.
Suppose that $R$ is not a Dedekind domain. By [11, Proposition 2.2], there exists a maximal ideal $M\subset R$ of height strictly bigger than one. If $M\cap\mathbb{Z}=(0)$, then, since $\mathbb{Z}[X]\subseteq R\subseteq\mathbb{Q}[X]$, it follows that $R_{\mathbb{Z}\setminus\{0\}}=\mathbb{Q}[X]$ and $12\leq htM=ht(M_{\mathbb{Z}\setminus\{0\}})\leq dim(\mathbb{Q}[X])=$ 1, a contradiction. Hence, $M\cap\mathbb{Z}=p\mathbb{Z}$, for some $p\in\mathbb{P}$. If we now localize at $p$, we have that $(\mathbb{Z}\setminus p\mathbb{Z})^{-1}R=R_{p}=\textnormal{Int}_{\mathbb{Q}}(E_{p},\mathbb{O}_{p})$ which is a Dedekind domain by [9, Theorem]. So $(\mathbb{Z}\setminus p\mathbb{Z})^{-1}M\subset R_{p}$ cannot have dimension strictly bigger than one, a contradiction.
@@
Finally, suppose that $R$ is a Dedekind domain. As in [19, Lemma 2.14], we have $\text{Cl}(R)=\bigoplus_{p\in\mathbb{P}}\text{Cl}(R_{p})$, where $R_{p}=\textnormal{Int}_{\mathbb{Q}}(E_{p},\mathbb{O}_{p})$ for $p\in\mathbb{P}$. The claim about the class group of $R_{p}$ follows by [19, Proposition 2.10] or by [9, Theorem], since, for each $p\in\mathbb{P}$, we are assuming that $E_{p}=\{\alpha_{1},\ldots,\alpha_{n}\}$ is formed by pairwise non-conjugate elements over $\mathbb{Q}_{p}$.
@@
Let ^¯$\overline{\widehat{\mathbb{Z}}}=\prod_{p\in\mathbb{P}}\overline{\mathbb{Z}_{p}}$. In [19, Theorem 2.17] we show that if $R$ is a Dedekind domain between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ such that the residue fields of prime characteristic are finite fields, then $R=\textnormal{Int}_{\mathbb{Q}}(\underline{E},\overline{\widehat{\mathbb{Z}}})$, for some $\underline{E}=\prod_{p}E_{p}\subset\overline{\widehat{\mathbb{Z}}}$ such that $\underline{E}$ is polynomially factorizable and for each $p\in\mathbb{P}$, $E_{p}$ is a finite subset of $\overline{\mathbb{Z}_{p}}$ of transcendental elements over $\mathbb{Q}$. Now, we are able to complete the classification of the Dedekind domains $R$, $\mathbb{Z}[X]\subset R\subseteq\mathbb{Q}[X]$, without any restriction on the residue fields.
@@
As we mentioned in the Introduction, given a group $G$ which is the direct sum $G$ of a countable family of finitely generated abelian groups, there exists a Dedekind domain $R$ between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ with class group $G$ ([19, Theorem 3.1]). The domain $R$ of that construction is equal to $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$ for some polynomially factorizable subset $\underline{E}=\prod_{p\in\mathbb{P}}E_{p}$, where $E_{p}$ is a finite subset of $\overline{\mathbb{Q}_{p}}$ of transcendental elements over $\mathbb{Q}$. In particular, $R$ has finite residue fields of prime characteristic ([19, Theorem 2.17]); the reason is that the valuation overrings $\mathbb{Z}_{(p),\alpha_{p}}$ of $R$ in (3.1) have finite residue fields precisely because $\alpha_{p}$ is chosen in $\overline{\mathbb{Q}_{p}}$ for each $p\in\mathbb{P}$ (Proposition 2.24).
Now, by means of Theorem 2.26, with the same method used in [19, Theorem 3.1], we can build a Dedekind domain $R$, $\mathbb{Z}[X]\subset R\subseteq\mathbb{Q}[X]$, with prescribed class group $G$ as above and prescribed residue fields of prime characteristic, which can be finite or infinite algebraic extensions of the prime field $\mathbb{F}_{p}$, according to whether the the above elements $\alpha_{p}\in E_{p}\subset\mathbb{C}_{p}^{\text{br}}$ transcendental over $\mathbb{Q}$, are either algebraic or transcendental over $\mathbb{Q}_{p}$.

### v3

@@ 2.3 Residually algebraic torsion extensions of $\mathbb{Z}_{(p)}$
The next proposition is analogous to Proposition 2.4, and characterizes residually algebraic torsion extensions of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$ in terms of pseudo-convergent sequences of $\overline{\mathbb{Q}_{p}}$ which are of transcendental type over $\mathbb{Q}$; clearly, every pseudo-convergent sequence of transcendental type of $\overline{\mathbb{Q}_{p}}$ belongs to this class. As a particular case, we find again part of the result of [19, Theorem 2.5].
@@
If $\widehat{\mathbb{Q}(X)}$ is a finite extension of $\mathbb{Q}_{p}$ then clearly $X$ is algebraic over $\mathbb{Q}_{p}$, so 1. implies 2. If $X$ is algebraic over $\mathbb{Q}_{p}$, we may identify $X$ with some $\alpha\in\overline{\mathbb{Q}_{p}}$; $\mathbb{Q}_{p}(\alpha)$ is a finite extension of $\mathbb{Q}_{p}$, hence complete. So, $\widehat{\mathbb{Q}(X)}=\mathbb{Q}_{p}(\alpha)$. As in the proof of [19, Theorem 2.5] it follows easily that $W=\mathbb{Z}_{(p),\alpha}$. Therefore, 2. implies 3.
If $W=\mathbb{Z}_{(p),\alpha}$ for some $\alpha\in\overline{\mathbb{Q}_{p}}$ transcendental over $\mathbb{Q}$, then by [19, Proposition 2.2], $ef<\infty$, so 3. implies 4. Finally, 4. implies 1. by [19, Lemma 2.4] because $e(\widehat{W}\mid\mathbb{Z}_{p})=e$ and $f(\widehat{W}\mid\mathbb{Z}_{p})=f$.
Note that if $E\subset\overline{\mathbb{Q}_{p}}$ is a pseudo-convergent sequence such that $\mathbb{Z}_{(p),E}=\mathbb{Z}_{(p),\alpha}$, then by Lemma 2.27 below we have $\mathbb{Z}_{p,E}=\mathbb{Z}_{p,\alpha}$, so by [23, Proposition 5.3] we have that $E$ is Cauchy and converges to $\alpha$.
The claims about the value group and residue field of $\mathbb{Z}_{(p),\alpha}$ follow by [19, Proposition 2.2].
@@
Let $W$ be a residually algebraic torsion extension of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$. Then the completion $\widehat{\mathbb{Q}(X)}$ with respect to $W$ is (isomorphic to) a subfield of $\mathbb{C}_{p}$ if and only if there exists $\alpha\in\mathbb{C}_{p}$, transcendental over $\mathbb{Q}$, such that $W=\mathbb{Z}_{(p),\alpha}$.
@@
According to Proposition 2.24, when passing to the completion, either $X$ is algebraic over $\mathbb{Q}_{p}$ or $X$ is transcendental over $\mathbb{Q}_{p}$ and consequently either $\widehat{\mathbb{Q}(X)}\subset\overline{\mathbb{Q}_{p}}\subset\mathbb{C}_{p}$ or $\mathbb{Q}_{p}(X)\subset\widehat{\mathbb{Q}(X)}$, respectively. In the first case, $W=\mathbb{Z}_{(p),\alpha}$ for some $\alpha\in\overline{\mathbb{Q}_{p}}\subset\mathbb{C}_{p}$ transcendental over $\mathbb{Q}$. In the second case, $\widehat{\mathbb{Q}_{p}(X)}=\widehat{\mathbb{Q}(X)}$, where the completion of $\mathbb{Q}_{p}(X)$ is considered with respect to the valuation domain $\widetilde{W}=\widehat{W}\cap\mathbb{Q}_{p}(X)$. In particular, by Proposition 2.23, we get that $\widehat{\mathbb{Q}(X)}\subseteq\mathbb{C}_{p}$ if and only if there exists a transcendental element $\alpha\in\mathbb{C}_{p}$ such that $W=\mathbb{Z}_{(p),\alpha}$. ∎
@@
The following result is the analogous of Theorem 2.21 for building residually algebraic torsion extensions $W$ of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$ with prescribed residue field $k$ and value group $\Gamma$. Note that, contrary to that Theorem, now we are not assuming anymore that $k:\mathbb{F}_{p}]\cdot[\Gamma:\mathbb{Z}]=\infty$.
@@
Let $e=[\Gamma:\mathbb{Z}]$ and $f=[k:\mathbb{F}_{p}]$. If $ef<\infty$, then it is well known that there exists $\alpha\in\overline{\mathbb{Q}_{p}}$ transcendental over $\mathbb{Q}$ such that $\mathbb{O}_{p,\alpha}$ has residue field $k$ and value group $\Gamma$. Hence, by [19, Proposition 2.2], $\mathbb{Z}_{(p),\alpha}$ is the desired extension of $\mathbb{Z}_{(p)}$.
@@
Therefore, if we choose $a_{i}\in\mathbb{Q}$ sufficiently $v_{p}$-adically close to $\alpha_{i}$ for each $i=0,\ldots,d$, we have $w(f-g)\geq 0$. In particular, $g\in W\cap\mathbb{Q}(X)=\mathbb{Z}_{(p),\alpha}$. The polynomial h=fgh=f-g is in $\mathbb{Z}_{p}[X]$; therefore $f(\alpha)=h(\alpha)+g(\alpha)\in\mathbb{O}_{p}$, so that $f\in\mathbb{Z}_{p,\alpha}$. Therefore $W\cap\mathbb{Q}_{p}[X]\subseteq\mathbb{Z}_{p,\alpha}\cap\mathbb{Q}_{p}[X]$. Similarly, one can easily show that the other containment holds, so $W\cap\mathbb{Q}_{p}[X]=\mathbb{Z}_{p,\alpha}\cap\mathbb{Q}_{p}[X]$. In the same way, $M_{W}\cap\mathbb{Q}_{p}[X]=M_{p,\alpha}\cap\mathbb{Q}_{p}[X]$, where $M_{W}$ and $M_{p,\alpha}$ are the maximal ideals of $W$ and $\mathbb{Z}_{p,\alpha}$, respectively.
@@
Let $f=[W/M:\mathbb{Z}/p\mathbb{Z}]$. If $f<\infty$, then the claim follows by [19, Theorem 2.5] and corresponds to the first case of Proposition 2.24: $W=\mathbb{Z}_{(p),\alpha}$, for some $\alpha\in\overline{\mathbb{Q}_{p}}$ which is transcendental over $\mathbb{Q}$. If $f=\infty$, then we are in the last case of Proposition 2.24, so $W=\mathbb{Z}_{(p),E}$, for some pseudo-convergent sequence in $\overline{\mathbb{Q}_{p}}$ of transcendental type. As in the proof of Proposition 2.24, we denote by $\widehat{W}$ the completion of $W$; since the ramification index $e(W\mid\mathbb{Z}_{(p)})$ is finite, $\widetilde{W}=\widehat{W}\cap\mathbb{Q}_{p}(X)$ is a residually algebraic torsion extension of $\mathbb{Z}_{p}$ to $\mathbb{Q}_{p}(X)$ which is a DVR, so by Theorem 2.17, $\widetilde{W}=\mathbb{Z}_{p,\alpha}$ for some $\alpha\in\mathbb{C}_{p}^{\text{br}}\setminus\overline{\mathbb{Q}_{p}}$. Hence, $W=\widetilde{W}\cap\mathbb{Q}(X)=\mathbb{Z}_{(p),\alpha}$. Note that $\alpha$ is transcendental over $\mathbb{Q}_{p}$, hence also over $\mathbb{Q}$.
@@
Conversely, suppose that $\mathbb{Z}_{(p),\alpha}=\mathbb{Z}_{(p),\beta}$. By Lemma 2.27, it follows that $\mathbb{Z}_{p,\alpha}=\mathbb{Z}_{p,\beta}$. Note that the last two valuation domains are the contraction to $\mathbb{Q}_{p}(X)$ of the valuation domains $\overline{\mathbb{Z}_{p}}_{\alpha}=\{\phi\in\overline{\mathbb{Q}_{p}}(X)\mid\phi(\alpha)\in\mathbb{O}_{p}\},\overline{\mathbb{Z}_{p}}_{\beta}=\{\phi\in\overline{\mathbb{Q}_{p}}(X)\mid\phi(\alpha)\in\mathbb{O}_{p}\}$ of $\overline{\mathbb{Q}_{p}}(X)$, respectively. By [6, Chapt. VI, §8, 6., Corollary 1], there exists a $\mathbb{Q}_{p}(X)$-automorphism $\sigma$ of $\overline{\mathbb{Q}_{p}}(X)$ such that $\sigma(\overline{\mathbb{Z}_{p}}_{\alpha})=\overline{\mathbb{Z}_{p}}_{\beta}$. It is easy to check that $\sigma(\overline{\mathbb{Z}_{p}}_{\alpha})=\overline{\mathbb{Z}_{p}}_{\sigma(\alpha)}$. In particular, $\overline{\mathbb{Z}_{p}}_{\sigma(\alpha)}=\overline{\mathbb{Z}_{p}}_{\beta}$. If $\sigma(\alpha)-\beta\not=0$, let $c\in\overline{\mathbb{Z}_{p}}$ be such that $v_{p}(c)>v_{p}(\sigma(\alpha)-\beta)$. Let $a\in\overline{\mathbb{Q}_{p}}$ be such that $v_{p}(a-\sigma(\alpha))\geq v_{p}(c)$. Then the polynomial $\frac{X-a}{c}$ is in $\overline{\mathbb{Z}_{p}}_{\sigma(\alpha)}$ and not in $\overline{\mathbb{Z}_{p}}_{\beta}$, a contradiction. ∎
@@ 3 Polynomial Dedekind domains
We remark that we have the following representation for the ring $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$ as an intersection of valuation overrings (see [21, (2.2)], for example):
@@
where $\mathcal{P}^{\text{irr}}$ denotes the set of irreducible polynomials in $\mathbb{Q}[X]$. By [19, Proposition 2.2], the valuation domain $\mathbb{Z}_{(p),\alpha_{p}}$ of $\mathbb{Q}(X)$ has rank one if and only if $\alpha_{p}$ is transcendental over $\mathbb{Q}$, and has rank $2$ otherwise (in the last case, note that necessarily $\alpha\in\overline{\mathbb{Q}_{p}}$).
A totally similar argument to [21, Lemma 2.5] shows that, for $p\in\mathbb{P}$, we have
@@
Suppose that the above conditions on $\underline{E}$ are satisfied. By (3.1), $R$ is equal to an intersection of DVRs. Moreover, $R$ has finite character, that is, for every non-zero $\in R, ff$ belongs to finitely many maximal ideals of the family of DVRs appearing in (3.1): in fact, if $f(X)=\frac{g(X)}{n}$, for some $g\in\mathbb{Z}[X]$ and $n\in\mathbb{Z}$, $n\not=0$, then $f$ is divisible only by finitely many $q\in\mathcal{P}^{\text{irr}}$; since $\underline{E}$ is polynomially factorizable, by [21, Lemma 2.12], the set $\{p\in\mathbb{P}\mid\exists\alpha_{p}\in E_{p},v_{p}(g(\alpha_{p}))>0\}$ is finite, so that $f$ belongs to finitely many maximal ideals of the family $\mathbb{Z}_{(p),\alpha_{p}}$, $\alpha_{p}\in E_{p},p\in\mathbb{P}$. Hence, $R$ is a Krull domain.
Suppose that $R$ is not a Dedekind domain. By [11, Proposition 2.2], there exists a maximal ideal $M\subset R$ of height strictly bigger than one. If $M\cap\mathbb{Z}=(0)$, then, since $\mathbb{Z}[X]\subseteq R\subseteq\mathbb{Q}[X]$, it follows that $R_{\mathbb{Z}\setminus\{0\}}=\mathbb{Q}[X]$ and $2\leq htM=ht(M_{\mathbb{Z}\setminus\{0\}})\leq dim(\mathbb{Q}[X])=1$, a contradiction. Hence, $M\cap\mathbb{Z}=p\mathbb{Z}$, for some $p\in\mathbb{P}$. If we now localize at $p$, we have that $(\mathbb{Z}\setminus p\mathbb{Z})^{-1}R=R_{p}=\textnormal{Int}_{\mathbb{Q}}(E_{p},\mathbb{O}_{p})$ which is a Dedekind domain by [9, Theorem]. So $(\mathbb{Z}\setminus p\mathbb{Z})^{-1}M\subset R_{p}$ cannot have dimension strictly bigger than one, a contradiction.
@@
Finally, suppose that $R$ is a Dedekind domain. As in [21, Lemma 2.14], we have $\text{Cl}(R)=\bigoplus_{p\in\mathbb{P}}\text{Cl}(R_{p})$, where $R_{p}=\textnormal{Int}_{\mathbb{Q}}(E_{p},\mathbb{O}_{p})$ for $p\in\mathbb{P}$. The claim about the class group of $R_{p}$ follows by [21, Proposition 2.10] or by [9, Theorem], since, for each $p\in\mathbb{P}$, we are assuming that $E_{p}=\{\alpha_{1},\ldots,\alpha_{n}\}$ is formed by pairwise non-conjugate elements over $\mathbb{Q}_{p}$.
@@
[…]
@@
As we mentioned in the Introduction, given a group $G$ which is the direct sum $G$ of a countable family of finitely generated abelian groups, there exists a Dedekind domain $R$ between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ with class group $G$ ([21, Theorem 3.1]). The domain $R$ of that construction is equal to $\textnormal{Int}_{\mathbb{Q}}(\underline{E},\mathcal{O})$ for some polynomially factorizable subset $\underline{E}=\prod_{p\in\mathbb{P}}E_{p}$, where $E_{p}$ is a finite subset of $\overline{\mathbb{Q}_{p}}$ of transcendental elements over $\mathbb{Q}$. In particular, $R$ has finite residue fields of prime characteristic ([21, Theorem 2.17]); the reason is that the valuation overrings $\mathbb{Z}_{(p),\alpha_{p}}$ of $R$ in (3.1) have finite residue fields precisely because $\alpha_{p}$ is chosen in $\overline{\mathbb{Q}_{p}}$ for each $p\in\mathbb{P}$ (Proposition 2.24).
Now, by means of Theorem 2.26, with the same method used in [21, Theorem 3.1], we can build a Dedekind domain $R$, $\mathbb{Z}[X]\subset R\subseteq\mathbb{Q}[X]$, with prescribed class group $G$ as above and prescribed residue fields of prime characteristic, which can be finite or infinite algebraic extensions of the prime field $\mathbb{F}_{p}$, according to whether the the above elements $\alpha_{p}\in E_{p}\subset\mathbb{C}_{p}^{\text{br}}$ transcendental over $\mathbb{Q}$, are either algebraic or transcendental over $\mathbb{Q}_{p}$.